OpenStudy (nannie):

Here is another. Please help! Divide and Simplify 11z-66/4÷z-6/6z

OpenStudy (anonymous):

=(11z-66)/4*6z/(z-6) =11(z-6)/4*6z/(z-60 =33z/2